OpenStudy (anonymous):

A length of a rectangle is four times the width. What is the perimeter of a rectangle with width 13 inches? A. 52 inches B. 60 inches C. 65 inches D. 130 inches

OpenStudy (anonymous):

help

OpenStudy (clalgee):

Hello @Crazy_questions ! I'll help you :)

OpenStudy (anonymous):

hello & thanks

OpenStudy (clalgee):

Alright, give me a few seconds. I type really slow :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

really call of duty

OpenStudy (clalgee):

Alright, So in the word sentence that we are given, it says, "the length of a rectangle is four times the width". We must write that into an equation. Equation: L = 4W Now lets put that aside for now. Now I recall that the perimeter of a rectangle is P=2L+2W. Right? Alright a few more seconds

OpenStudy (anonymous):

brb

OpenStudy (clalgee):

Now this is our new equation: P=2(4W)+2W = 10W Now we have to replace the W with 13. What does that equal? Just multiply 13 x 10 and you have your answer! :)

OpenStudy (anonymous):

back

OpenStudy (clalgee):

Do you think you know the answer now? (13 x 10 = ? )

OpenStudy (anonymous):

13*10=130?

OpenStudy (clalgee):

Yep!

OpenStudy (clalgee):

So, looking at the choices given. Which one is it? :)

OpenStudy (anonymous):

ok thank you for the help and its D

OpenStudy (clalgee):

Yes :D ! And your welcome! Let me know if you need any more help and I will be there :)

OpenStudy (anonymous):

ok thanks

OpenStudy (clalgee):

Np :)

OpenStudy (anonymous):

bye :-)